Prev Next

SW&IT Meeting Minutes 2020-02-17

Attendees: Carl, David, Igor, Matthieu, Petr, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200210
LR20200210 - Bec, Matthieu 40.0
LR20200210 - Holmberg, Carl 32.0
LR20200210 - Edmeades, David 40.0
LR20200210 - Sola, Igor 40.0
LR20200210 - Kubanek, Petr 28.0
LR20200210 - Hooper, Stephen 80.0
LR20200210 - Zhang, Yang 40.0
LR20200210 - Zhang, Xianyu 56.0

News from management

Previous Action Items

OwnerSorted ascending Action Due Date Completion Date
DE Summit Veam Backup (soul-telemetry, etc) 2020-01-20 2020-02-14
DE timecapsule obsdev 2020-02-12  
DE oVirt for PEPSI 2020-02-12  
IS Luci passwords in google doc 2020-02-10  

Open topics

Team reports

YZ

  • MODS file checksum (cron task running + OSU started their work), created "lbc-alerts@lbto.org" and "mods-alerts@lbto.org" for LBC and MODS related alarms.
  • Updating LBC daily tech image moving script.
  • pmac patch and M2 firmware (DX adsec & rigid) support on Saturday & Sunday.
  • TMS software meeting on Friday.
  • Need to schedule some time with Olga for scripts on 64-bit machine.
  • Investigate ECS instrument cooling GUI color scheme.
  • AO-Cravat restart.

SH

  • Weather underground cameras for MGIO
  • KVM back up the mountain (diagnosed as power supply, David purchased another).
  • Updated NUC software.
  • obs1 Xorg, rm507 stabilization
  • Telemetry script to push to Tucson (fixed on RS2).

DE

  • DT tape hoard disposal
    • Contacted recycle company, days later I have nothing. Thinking that we just landfill them
  • Summit laptop feature freeze at OT and wiki
  • Conversation with Veeam satisfied my questions, ordered license for mountain VM host.
  • UITS email hiccup Thursday evening
    • Spam filter blocked a message from a Google domain and decided to blacklist all Google domains including LBTO
    • UITS located and fixed the problem
  • 4k obs display kick?

MB

  • Last week
  • This week
    • ROWG kickoff
    • TMS #2
    • TCS ITs review with PK and YZ
    • DMS developments

PK

  • pmac patch, making sure firmware loads properly
  • GCS patch for Zernikes index
  • DX M2 hexapod help
  • IT 5140 - discovered issue with pcs.monocularMode
  • preparing presentation for UMAC training

CH

Last week

  • AHS: S. Ertel provided feedback on LBTI alert approach. I think his answers are sufficient for us to proceed.
  • ARGOS: Reviewed code for data entry points.
  • Support scripts: tested updated RB_Science on robs.mountain.lbto.org with work files moved to local fs, test files with FITSVERIFY. Still bails after ~5 image pairs. Perhaps update pyds9 with MB's newer RPM?

This week

  • AHS: Emplement LBTI alert script, update AHS to accept the new input.
  • ARGOS: Reviewed code for data entry points.

IS

Last week:
  • PIT
    • Migration of bug fixes to Italian PIT
This week:
  • OT
    • Final release?
  • PIT
    • PIT enhancements

  • Time Accounting Tool?
    • Export to excel
    • Use django admin login capabilities for login and queue run program editing.
Topic revision: r12 - 17 Feb 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback